Suppose that a 1-year zero-coupon bond with face value $100 currently sells at $90.34, while a 2-year zero sells at $81.43. You are considering the purchase of a 2-year-maturity bond making annual coupon payments. The face value of the bond is $100, and the coupon rate is 16% per year.

a. What is the yield to maturity of the 2-year zero?(Do not round intermediate calculations. Round your answers to 3 decimal places.)

b. What is the yield to maturity of the 2-year coupon bond? (Do not round intermediate calculations. Round your answers to 3 decimal places.)


c. What is the forward rate for the second year? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)


d. If the expectations hypothesis is accepted, what are (1) the expected price of the coupon bond at the end of the first year and (2) the expected holding-period return on the coupon bond over the first year? (Do not round intermediate calculations. Round your answers to 2 decimal places.)

e. Will the expected rate of return be higher or lower if you accept the liquidity preference hypothesis?

  • Higher

  • Lower
Answer #1

1.
=(100/81.43)^(1/2)-1=10.8173556229247%

2.
=RATE(2,16%*100,-16%*100*0.9034-(100+16%*100)*0.8143,100)=10.8085131152888%

3.
=90.34/81.43-1=10.9419%

4.
Expected price=(100+16%*100)/(1+10.9419%)=104.5592

Expected holding period return=(104.5592+16%*100)/(16%*100*0.9034+(100+16%*100)*0.8143)-1=10.6929%

5.
Lower